• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name kamyers

    Log entry time 10:03:27 on March30,2012

    Entry number 368341

    keyword=Beam current 50 nA -> 65 nA

    During runs 3497/22531 we increased the beam current from 50 nA to 65 nA to bring the DAQ rate up (we had ~8% DT and were stuck at the prescale setting). The beam positions at IPM1H05B drifted by ~0.5 mm, and the A bpm seems more noisey? The target polarization is dropping more quickly, so the RC is debating if we want to continue running this way

    FIGURE 1